sphere of mass M is held between 2 smooth inclined planes for Sin 37 is equal to 3 by 5 the normal reaction of the wall 2 to is equal to

Grade:

1 Answers

bramashini
5 years ago
Mg=2Rsin37
R=Mg/2sin37
R=Mg / 2*3/5    ( sin 37=3/5 is given)
 therefore , R=5Mg/6
this sphere is lying  between  2 inclined planes so there are 2 normal reactions on the sphere by the plane. if we resolve these reactions the vertical forces will be Rsin37*2 this system is under  equllibrium so verical forces 2Rsin37 is equal to weight of the sphere Mg (here assume normal reaction= R)